OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 8 PI A= ? PI

OpenStudy (anonymous):

cirumference of a cirlce is P = 2*pi*r, so solve for r which gives r=4units then area = pi*r^2 = 3.14*(4)^2 = 50.24sq.units. bamf

OpenStudy (anonymous):

but i'm not finding square units anymore aha its finding the exact area of PI

OpenStudy (anonymous):

ahh well, then it would just be 16pi